OpenStudy (anonymous):

Find the sum of the geometric series: 8 + 4 + 2 + 1 + ...

OpenStudy (ybarrap):

down to zero?

OpenStudy (anonymous):

umm it doesn't say it just says....

OpenStudy (ybarrap):

\[\sum_{0}^{3}2^{n}= (1 - 2^{4})/(1- 2)\]

OpenStudy (anonymous):

\[r=\frac{ 4 }{ 8}=\frac{ 1 }{ 2 }<1,a=8\] \[S \infty=\frac{ a }{ 1-r }\] substitute the value of a and r get the value of S infty

OpenStudy (nurali):

a = 8 r = 1/2 S = a/(1-r) Note: this formula only works if |r| < 1, which is true in this case (since |1/2| < 1 is true) S=16

OpenStudy (anonymous):

ok thanks guys:)

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

how abt this one?

OpenStudy (ybarrap):

17/.7

OpenStudy (anonymous):

why would it be over .7?

OpenStudy (ybarrap):

\[\sum_{i=1}^{\infty}0.3^{i} = 1/(1-.3)\]

OpenStudy (ybarrap):

That should have been \sum_{i=0}^{\infty}0.3^{i} = 1/(1-.3)

OpenStudy (ybarrap):

\[sum_{i=1}^{\infty}0.3^{i} = 1/(1-.3)\]

OpenStudy (anonymous):

then what abt the 17???

OpenStudy (ybarrap):

That's just a factor multiplied times the whole sum, nothing special.

OpenStudy (ybarrap):

so multiply 1/.7 times 17

OpenStudy (ybarrap):

like this, 17(.3^0 + .3^1 + .3^2 + ...)

OpenStudy (anonymous):

so 170/7?

OpenStudy (ybarrap):

yes, same thing, that's just 17/.7 * 1 = 17/.7 * 10/10

OpenStudy (ybarrap):

for those that don't like mixing decimals and fractions

OpenStudy (anonymous):

ok thnks :)

OpenStudy (ybarrap):

not a problem